1.      Consider in perfectly competitive market the following demand and supply equations for sugar:

Qd =1000-1000p        where Q d is quantity demanded and Qs is quantity supplied. Qs=800+ 1000p

Where P is the price of sugar per pound and Q is thousands of pounds of sugar.


1
Expert's answer
2021-04-22T09:38:36-0400

In equilibrium Qd = Qs, so:

1000 - 1000p = 800 + 1000p,

2000p = 200,

p = 0.1,

Q = 1000 - 1000×0.1 = 900 thousands of pounds of sugar.